calculate the ss for the following sample of n=8 scores

0,4,1,3,2,1,1,0


1
Expert's answer
2022-02-07T17:40:32-0500

Solution;

"\begin{array}{cc}
 x & 0&4&1&3&2&1&1&0&\sum x=12 \\
 x^2 & 0&16&1&9&4&1&1&0& \sum x^2=32
\end{array}"

Therefore;

The sum of squares,SS;

"SS=\sum x^2-\frac{(\sum x)^2}{n}"

"SS=32-\frac{12}{8}^2=14"